Assume that you plan to buy a condo 5 years from now, and you estimate that you can save $3 ,500 per year. You plan to deposit the money in a bank that pays 4% interest, and you will make the first deposit at the end of the year. How much will you have after 5 years? How will your answer change if the interest rate is increased to 6% or lowered to 3%?
Answer #1

N = 5 years

PMT = 3,500

I = 4%

FV = 3,500 * (1 + 4%)4 + 3,500 * (1 + 4%)3 + 3,500 * (1 + 4%)2 + 3,500 * (1 + 4%) + 3,500

FV = 4,094.50 + 3,937.02 + 3,785.60 + 3,640 + 3,500

FV = 18,957.13

Part 2:

N = 5 years

PMT = 3,500

I = 6%

FV = 3,500 * (1 + 6%)4 + 3,500 * (1 + 6%)3 + 3,500 * (1 + 6%)2 + 3,500 * (1 + 6%) + 3,500

FV = 4,418.67 + 4,168.56 + 3,932.60 + 3,710 + 3,500

FV = 19,729.83

Part 3:

N = 5 years

PMT = 3,500

I = 3%

FV = 3,500 * (1 + 3%)4 + 3,500 * (1 + 3%)3 + 3,500 * (1 + 3%)2 + 3,500 * (1 + 3%) + 3,500

FV = 3,939.28 + 3,824.54 + 3,713.15 + 3,605 + 3,500

FV = 18,581.98